in ten seconds, what will be the distance that sound can travel through air at 50⁰C?
1
Expert's answer
2020-11-05T03:53:37-0500

The equation for the speed of sound in air is the following.


See https://phys.libretexts.org/Bookshelves/University_Physics/Book%3A_University_Physics_(OpenStax)/Map%3A_University_Physics_I_-_Mechanics_Sound_Oscillations_and_Waves_(OpenStax)/17%3A_Sound/17.03%3A_Speed_of_Sound#:~:text=If%20the%20temperature%20is%20T,s√T273K.&text=v%3Dfλ



"v = 331m/s\sqrt{\dfrac{T}{273K}}"

where "T = 273K+50\degree C = 323K" is the absolute temperature of air. Thus, obtain:


"v = 331m/s\sqrt{\dfrac{323K}{273K}} \approx 360 m/s"

Answer. 360 m/s.